# Practical hybrid PQC-QKD protocols with enhanced security and performance

### v2

@@ Abstract
Quantum resistance is vital for emerging cryptographic systems as quantum technologies continue to advance towards large-scale, fault-tolerant quantum computers. Resistance may be offered by quantum key distribution (QKD), which provides information-theoretic security using quantum states of photons, but may be limited by transmission loss at long distances. An alternative approach uses classical means and is conjectured to be resistant to quantum attacks—so-called post-quantum cryptography (PQC)—but it is yet to be rigorously proven, and its current implementations are computationally expensive. To overcome the security and performance challenges present in each, here we develop hybrid protocols by which QKD and PQC inter-operate within a joint quantum-classical network. In particular, we consider different hybrid designs that may offer enhanced speed and/or security over the individual performance of either approach. Furthermore, we present a method for analyzing the security of hybrid protocols in key distribution networks. Our hybrid approach paves the way for joint quantum-classical communication networks, which leverage the advantages of both QKD and PQC and can be tailored to the requirements of various practical networks.
@@
Key Encapsulation Mechanism and Quantum Key Distribution. One major component we use to construct the symmetric key distribution network is prepare-and measure QKD [3, 11], shown in Fig. 1(a). In this protocol Alice generates a random raw key $rk$, which she encodes into quantum states that she transmits to Bob. She records classical information $b$ about the encoding such as the basis used. Bob receives the quantum states and measures them to obtain $rk^{\prime}$ and classical information $b^{\prime}$ related to the measurement procedure. Alice and Bob then announce $b$ and $b^{\prime}$ as well as a subset of $rk$ and $rk^{\prime}$ and perform classical post-processing, including security parameter estimation, error correction, and privacy amplification, to generate the final symmetric key bits $k$ from the raw key bits $rk(rk^{\prime})$.
@@
Security enhancement by parallel design. Another motivation for combining KEM and QKD is to achieve a higher level of security. Both bare KEM and QKD protocols have security vulnerabilities when deployed in practice: KEM may eventually be broken algorithmically, posing a risk of becoming unreliable in the future, while current implementations of QKD may be susceptible to physical attacks due to device imperfections [11]. Additionally, in scenarios where a QKD link involves classical relay nodes or the series-connection protocol discussed earlier, it becomes necessary to assume the trustworthiness of all intermediate nodes. In the series-connection protocol, if Eve manages to compromise either the QKD or KEM link, she could potentially access the final key bits shared by the users. To address this, we explore how end users can enhance the security of distributed key bits when utilizing multiple key distribution channels. Specifically, we propose two parallel key-distribution designs: the XOR scheme and the secret-sharing (SS) scheme.
@@
Network analysis. In a realistic hybrid quantum network, information may need to travel through intermediary links and nodes before it reaches its destination. To assess the security of a protocol implemented in such a network one must consider the vulnerability of all the involved links and nodes. For example, consider replacing a long QKD link between Alice and Bob with many short QKD links in series with intermediate trusted nodes. Since the links are shorter, the overall key rate is higher, but the trusted nodes can increase the vulnerability of the protocol. In such a scheme, if even one of the nodes is compromised, then the secret key can be fully exposed. Vyas and Mendes [42] suggest a protocol to relax the trust requirement by connecting each node to a highly trusted central key management system (KMS). Instead of using the generated QKD keys to decrypt and re-encrypt information, each node now takes the XOR of the keys distributed over its two adjacent links to generate a mask, which is then submitted to the KMS. In this scheme, no single node except for Alice and Bob ever has enough information to learn the secret key, assuming communication with the KMS is secured (for example, through KEM). To learn the secret key, an attacker would need to learn not only a raw QKD key, but also all of the masks (by compromising the KMS node or its communication links).
More generally, the security of hybrid protocols that build upon the series and parallel combination of QKD and KEM can be examined by representing a hybrid network as a graph $\mathcal{G}=(\mathcal{N},\mathcal{E})$, where $\mathcal{N}$ is the set of nodes and $\mathcal{E}$ is the set of QKD and KEM links between the nodes. We allow for $\mathcal{G}$ to have multiple edges between the same pair of nodes. This corresponds, for example, to the presence of both a QKD link and a KEM link between two nodes. A key sharing protocol $P_{A,B}$ between nodes $A$ and $B$ can be constructed from sub-protocols combined in series or in parallel, using the techniques described in previous sections. This protocol can be compromised by attacking some subset of the edges or nodes it uses. Formally, a vulnerability of a protocol is a set $v\in 2^{\mathcal{N}\cup\mathcal{E}}$ of network elements that, if attacked as a unit, would expose the shared key. We define the total vulnerability set of a protocol $V_{\text{tot}}(P_{A,B})$ as the set containing all possible vulnerabilities for protocol $P_{A,B}$. The minimal vulnerability set $V_{\text{min}}(P_{A,B})$ is the subset of $V_{\text{tot}}(P_{A,B})$ containing all of the vulnerabilities that have no strict subset in $V_{\text{tot}}(P_{A,B})$, thus describing the smallest set containing all the units that Eve could choose to attack in order to get the final key.
@@ Acknowledgements.
We thank Jens Eisert, Yi Li, Mouktik Raha, Grant Smith, Han Zheng, and Changchun Zhong for fruitful discussions. P.Z., D.B., and L.J. acknowledge support from the ARO(W911NF-23-1-0077), ARO MURI (W911NF-21-1-0325), AFOSR MURI (FA9550-19-1-0399, FA9550-21-1-0209, FA9550-23-1-0338), NSF (OMA-1936118, ERC-1941583, OMA-2137642, OSI-2326767, CCF-2312755), Packard Foundation (2020-71479), and the Marshall and Arlene Bennett Family Research Program. F.R. likewise acknowledges support from the NSF CQN (ERC-1941583). P.Z., D.B., and L.J. further acknowledge that this material is based upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers and Advanced Scientific Computing Research (ASCR) program under contract number DE-AC02-06CH11357 as part of the InterQnet quantum networking project. J.A.M.M., N.B., A.K., M.T.S., T.Z., F.J.H., and D.D.A. acknowledge additional support provided by Q-NEXT, part of the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, and the AFOSR MURI (FA9550-23-1-0330). J.L. acknowledges startup funds provided by the University of Pittsburgh and funding from IBM Quantum through the Chicago Quantum Exchange.

### v3

@@ Abstract
Quantum resistance is vital for emerging cryptographic systems as quantum technologies continue to advance towards large-scale, fault-tolerant quantum computers. Resistance may be offered by quantum key distribution (QKD), which provides information-theoretic security using quantum states of photons, but may be limited by transmission loss at long distances. An alternative approach uses classical means and is conjectured to be resistant to quantum attacks—so-called post-quantum cryptography (PQC)—but it is yet to be rigorously proven, and its current implementations are computationally expensive. To overcome the security and performance challenges present in each, here we develop hybrid protocols by which QKD and PQC inter-operate within a joint quantum-classical network. In particular, we consider different hybrid designs that may offer enhanced speed and/or security over the individual performance of either approach. Furthermore, we present a method for analyzing the security and performance of hybrid protocols in key distribution networks. Our hybrid approach paves the way for joint quantum-classical communication networks, which leverage the advantages of both QKD and PQC and can be tailored to the requirements of various practical networks.
@@
One major component we use to construct the symmetric key distribution network is prepare-and measure QKD [3, 11], shown in Fig. 1(a). In this protocol Alice generates a random raw key $rk$, which she encodes into quantum states that she transmits to Bob. She records classical information $b$ about the encoding such as the basis used. Bob receives the quantum states and measures them to obtain $rk^{\prime}$ and classical information $b^{\prime}$ related to the measurement procedure. Alice and Bob then announce $b$ and $b^{\prime}$ as well as a subset of $rk$ and $rk^{\prime}$ and perform classical post-processing, including security parameter estimation, error correction, and privacy amplification, to generate the final symmetric key bits $k$ from the raw key bits $rk(rk^{\prime})$.
@@
Another motivation for combining KEM and QKD is to achieve a higher level of security. Both bare KEM and QKD protocols have security vulnerabilities when deployed in practice: KEM may eventually be broken algorithmically, posing a risk of becoming unreliable in the future, while current implementations of QKD may be susceptible to physical attacks due to device imperfections [11]. Additionally, in scenarios where a QKD link involves classical relay nodes or the series-connection protocol discussed earlier, it becomes necessary to assume the trustworthiness of all intermediate nodes. In the series-connection protocol, if Eve manages to compromise either the QKD or KEM link, she could potentially access the final key bits shared by the users. To address this, we explore how end users can enhance the security of distributed key bits when utilizing multiple key distribution channels. Specifically, we propose two parallel key-distribution designs: the XOR scheme and the secret-sharing (SS) scheme.
@@
In a realistic hybrid quantum network, information may need to travel through intermediary links and nodes before it reaches its destination. To assess the security of a protocol implemented in such a network one must consider the vulnerability of all the involved links and nodes. For example, consider replacing a long QKD link between Alice and Bob with many short QKD links in series with intermediate trusted nodes. Since the links are shorter, the overall key rate is higher, but the trusted nodes can increase the vulnerability of the protocol. In such a scheme, if even one of the nodes is compromised, then the secret key can be fully exposed. Vyas and Mendes [42] suggest a protocol to relax the trust requirement by connecting each node to a highly trusted central key management system (KMS). Instead of using the generated QKD keys to decrypt and re-encrypt information, each node now takes the XOR of the keys distributed over its two adjacent links to generate a mask, which is then submitted to the KMS. In this scheme, no single node except for Alice and Bob ever has enough information to learn the secret key, assuming communication with the KMS is secured (for example, through KEM). To learn the secret key, an attacker would need to learn not only a raw QKD key, but also all of the masks (by compromising the KMS node or its communication links).
More generally, the security of hybrid protocols that build upon the series and parallel combination of QKD and KEM can be examined by representing a hybrid network as a graph $\mathcal{G}=(\mathcal{N},\mathcal{E})$, where $\mathcal{N}$ is the set of nodes and $\mathcal{E}$ is the set of QKD and KEM links between the nodes. We allow for $\mathcal{G}$ to have multiple edges between the same pair of nodes. This corresponds, for example, to the presence of both a QKD link and a KEM link between two nodes. A key sharing protocol $P_{A,B}$ between nodes $A$ and $B$ can be constructed from sub-protocols combined in series or in parallel, using the techniques described earlier. This protocol can be compromised by attacking some subset of the edges or nodes it uses. Formally, a vulnerability of a protocol is a set $v\in 2^{\mathcal{N}\cup\mathcal{E}}$ of network elements that, if attacked as a unit, would expose the shared key. We define the total vulnerability set of a protocol $V_{\text{tot}}(P_{A,B})$ as the set containing all possible vulnerabilities for protocol $P_{A,B}$. The minimal vulnerability set $V_{\text{min}}(P_{A,B})$ is the subset of $V_{\text{tot}}(P_{A,B})$ containing all of the vulnerabilities that have no strict subset in $V_{\text{tot}}(P_{A,B})$, thus describing the smallest set containing all the units that Eve could choose to attack in order to get the final key.
@@ Acknowledgements.
We thank Jens Eisert, Yi Li, Mouktik Raha, Grant Smith, Han Zheng, and Changchun Zhong for fruitful discussions. P.Z., D.B., and L.J. acknowledge support from the ARO (W911NF-23-1-0077), ARO MURI (W911NF-21-1-0325), AFOSR MURI (FA9550-19-1-0399, FA9550-21-1-0209, FA9550-23-1-0338), DARPA (HR0011-24-9-0359, HR0011-24-9-0361), NSF (OMA-1936118, ERC-1941583, OMA-2137642, OSI-2326767, CCF-2312755), NTT Research, Packard Foundation (2020-71479), and the Marshall and Arlene Bennett Family Research Program. F.R. likewise acknowledges support from the NSF CQN (ERC-1941583). P.Z., D.B., and L.J. further acknowledge that this material is based upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers and Advanced Scientific Computing Research (ASCR) program under contract number DE-AC02-06CH11357 as part of the InterQnet quantum networking project. J.A.M.M., N.B., A.K., M.T.S., T.Z., F.J.H., and D.D.A. acknowledge additional support provided by Q-NEXT, part of the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, and the AFOSR MURI (FA9550-23-1-0330). J.L. acknowledges startup funds provided by the University of Pittsburgh and funding from IBM Quantum through the Chicago Quantum Exchange.